# DrivAerNet: A Parametric Car Dataset for Data-Driven Aerodynamic Design and Graph-Based Drag Prediction Аннотация

### v1

@@ 3 DrivAerNet Dataset
DrivAer model [18] is a well-established conventional car reference model developed by researchers at the Technical University of Munich (TUM). It is a combination of the BMW series 3 and Audi A4 car designs in order to be representative of most conventional cars. The DrivAer model was developed to bridge the gap between open-source oversimplified models like the Ahmed [2] and SAE [11] bodies and the complex designs of the manufacturing companies, which are not publicly available. To accurately assess real-world aerodynamic designs, we selected the fastback configuration with a detailed underbody, wheels, and mirrors (FDwWwM) as our baseline model, as shown in Figure 1(a). This choice of the FDwWwM model was driven by the substantial impact of wheels, mirrors, and underbody geometry on aerodynamic drag, a conclusion supported by the findings in [17]. Specifically, the detailed underbody geometry adds 32-34 counts to the drag, the inclusion of mirrors introduces an additional 14-16 counts, and the presence of wheels elevates the total drag coefficient by 102 counts 222In aerodynamic evaluations, “drag counts” are used as a unit of measurement to denote small changes in the drag coefficient (). One drag count is defined as a increment in ..
@@
In order to create a comprehensive dataset for training deep learning models for surrogate modeling and design optimization, we first created a parametric model of the DrivAer model. This approach was necessitated by the limitations of the original model, which was provided as a single, non-parametric .stl file. To adequately capture the geometric variations and design modifications relevant to real-world automotive design challenges, we developed a version of the DrivAer model that is defined by 50 geometric parameters and includes 32 morphable entities (see Figure 1) using the commercial software ANSA®. This parametric model allows for a more detailed exploration of the design space, facilitating the generation of 4,000 unique design variants through the application of the Optimal Latin Hypercube sampling method, specifically employing the Enhanced Stochastic Evolutionary Algorithm (ESE) as outlined by [12].
@@
The parametric model, along with the constraints and the bounds applied during the Design of Experiments (DoE), significantly enriches the dataset, making it a robust foundation for the development and training of advanced deep learning models aimed at surrogate modeling and design optimization tasks. We provide access to the parametric model with the incorporated morphing features for further reference and utilization333The parametric model is available at https://github.com/Mohamedelrefaie/DrivAerNet/ParametricModel.
Diverging from the approach taken in [19], we implemented a broader range of morphing techniques, enabling us to explore a more diverse array of car designs. This approach aims to enhance the adaptability of deep learning models, allowing them to generalize across various car designs instead of being limited to minor geometric modifications within a single design. Figure 3 depicts the variation in mesh quality, ranging from coarse to high-resolution across various datasets. Our original mesh with 540k mesh faces, provides a denser and more detailed representation compared to the mesh resolutions in the studies by [22] and [36], thereby revealing more detailed geometric and design features.
@@ 3.1.1 Domain and Boundary Conditions
The DrivAer fastback model scaled 1:1, was selected for conducting the CFD simulations. The simulations were carried out using the open-source software OpenFOAM®, a comprehensive collection of C++ modules for tailoring solvers and utilities in CFD studies. In this study, the coupling between pressure and velocity is achieved through the SIMPLE algorithm (Semi-Implicit Method for Pressure Linked Equations), as implemented in the simpleFoam solver, which is designed for steady-state, turbulent, and incompressible flow simulations. The --SST model, based on Menter’s formulation [24], was chosen for the Reynolds-Averaged Navier-Stokes (RANS) simulations due to its ability to overcome the limitations of the standard - model, particularly its dependency on the freestream values of and , and its effectiveness in predicting flow separation.
The simulations were performed at a flow velocity () of 30 m/s, which corresponds to a Reynolds number of roughly , using the car length as the characteristic length scale. The computational mesh is constructed using the SnappyHexMesh (SHM) tool, featuring four distinct refinement zones. Additional layers have been added around the car body to precisely represent wake dynamics and boundary layer evolution (see Figure 1(b)). Boundary conditions were defined with a uniform velocity at the inlet and pressure-based conditions at the outlet. To avoid backflow into the simulation domain, the velocity boundary condition at the outlet was configured as an inletOutlet condition. The car surface and ground were assigned no-slip conditions, while the wheels were modeled to rotate with the rotatingWallVelocity boundary condition. Slip conditions were applied to the lateral and top boundaries of the domain.
@@ 3.1.2 Validation of the Numerical Results
| RANS medium | 205 | 0.81 | 1.88 | ||
| RANS fine | 512 | 5.7 | 6.78 | ||
The selection of the DrivAer fastback model is justified by the availability of both computational and experimental references, enabling us to benchmark our results against established data [17, 43]. Before commencing the simulations, we conducted a preliminary assessment of how mesh refinement influences the results. This involved comparing the drag coefficients obtained from three different mesh resolutions with experimental values and reference simulation, as detailed in Table 2. The objective was to identify an optimal balance between simulation accuracy and computational efficiency. This balance is crucial since our aim is to generate a large-scale dataset for training deep learning models, necessitating both high fidelity in the simulation results and manageable disk storage and simulation times to accommodate the extensive computational requirements. The drag coefficient, , is determined by the equation:
@@
Our analysis reveals a consistent correlation between our simulations and both the benchmark experimental data and reference simulations, with the 8 and 16 million cell meshes showing particularly good alignment. The discrepancy observed in the 40 million cell mesh could stem from differences in mesh granularity, as the reference simulations utilized a 16 million cell mesh. Finer meshes capture more intricate flow dynamics, which may not be represented in coarser meshes, leading to divergences. In computational fluid dynamics, especially in the preliminary stages of design, an error margin of up to 5 is generally considered acceptable for engineering purposes. In addition, implementing "RANS fine"for all 4000 designs within the DrivAerNet dataset would require approximately 120TB of storage, posing significant challenges to data sharing and reproducibility due to the immense storage demands. Therefore, considering the balance between accuracy and computational resource allocation, we decided to conduct our simulations using the 8 million and 16 million cell meshes. These configurations provide a compromise between computational efficiency and the level of detail necessary for accurate aerodynamic analysis.
Leveraging Multi-Fidelity Data and Transfer Learning for Efficient Surrogate Model Development:
As shown by [15], leveraging multi-fidelity CFD simulations proves to be a robust strategy for accurate 3D flow field estimation. This approach involves using a dataset that combines RANS, which are relatively easier and cheaper to obtain and can capture the general flow behavior, with Direct Numerical Simulation (DNS) data, known for its detailed flow information despite being computationally expensive. Training a deep learning model with this diverse dataset not only enables the model to generalize effectively to real-world scenarios, as confirmed by wind tunnel tests, but also facilitates a two-phase training process. This process starts with medium-fidelity RANS data to grasp general flow patterns and then transitions to fine-tuning with high-fidelity DNS data, thereby enhancing the model’s accuracy and real-world applicability. Similar results leveraging multi-fidelity datasets for training surrogate models have been demonstrated by [33, 35], underscoring the approach’s efficacy in aerodynamic analysis. The DrivAerNet dataset can be similarly utilized, allowing for integration with datasets of either lower or higher fidelity to augment model training and refine predictive capabilities.
@@ 3.1.3 CFD Simulation Results
In contrast to the approach by [36] where all car models are standardized to a uniform length of 3.5 meters to fit a predefined computational domain, our dataset allows diversity in car dimensions, adjusting the mesh, bounding boxes, and additional layers accordingly for each design. This flexibility is crucial for capturing the intricate flow dynamics around cars including phenomena like flow separation, reattachment, and recirculation zones, and ensuring precise aerodynamic coefficient estimation. This approach addresses limitations observed in some studies that prioritize dataset size over simulation fidelity, often overlooking the importance of convergence, accurate modeling, and appropriate boundary conditions for complex 3D models.
Modeling of Wheels, Side Mirrors, and the Underbody:
As previously highlighted, the majority of literature and available datasets tend to neglect the modeling of wheels, side mirrors, and the underbody, as detailed in Table 1. Our methodology, in contrast, includes detailed modeling of these components. Figure 5 illustrates the velocity distribution on the car: here, the car body exhibits zero velocity due to the no-slip boundary condition, whereas the wheels show nonzero velocity. Additionally, the figure visualizes streamlines around the car, offering insight into the flow dynamics influenced by the inclusion of these features. The DrivAerNet dataset features full 3D flow field information, as shown in Figure 5(a) with the velocity data, and in addition, it provides the pressure distribution on the car’s surface. The pressure coefficient, denoted as , is calculated as the ratio of the pressure differential to the dynamic pressure, , and is given by:
The distribution of on the car surface is depicted in Figure 5(b).
@@ 3.2 Geometric Feasibility
In our approach to generating large diverse automotive designs through automated morphing with ANSA®, ensuring the geometric quality and feasibility of each variant is crucial. To address potential issues such as non-watertight geometries, surface intersections, or internal holes resulting from morphing operations, we employ an automated mesh quality assessment and repair process. This procedure not only identifies but also rectifies common geometric anomalies, ensuring that only simulation-ready models are included in our dataset. Geometries failing to meet these criteria are systematically excluded from simulations. The DrivAerNet dataset employs a diverse array of parameters (total of 50 parameters) to morph car geometries, encompassing modifications to the side mirror placement, muffler position, windscreen, rear window length/inclination, the size of the engine undercover, offsets for the doors and fenders, hood positioning, and the scale of headlights as well as alterations to the overall car length and width. Additionally, adjustments are made to the car’s upper and underbody scaling, as well as pivotal angles such as the ramp, diffusor, and trunk lid angles, all crucial for exploring the effects of different design modifications on car aerodynamics. For a detailed account of the morphing parameters, including their lower and upper bounds, please refer to our GitHub repository.
As we morph the entire car’s geometry, wheel positioning is adjusted in the , , and axes during the morphing process. For all simulations, we use front and rear wheels of the same shape. To accurately simulate the wheel rotation, we export them as separate .stl files post-morphing, which allows us to apply the rotatingWallVelocity physical boundary condition. Moreover, morphing affects the car’s vertical positioning, necessitating a calculation of the -axis displacement to ensure the car body and wheels are properly aligned with the ground plane. For simulation purposes, we supply three distinct .stls: one for the car body, one for the front wheels, and one for the rear wheels, to model their interactions accurately.
@@ 3.3 DrivAerNet Dataset Characteristics
For our simulations, we employed OpenFOAM® version 11, executing the computational tasks across 128 CPU cores and 4 Nvidia A100 80GB GPUs. This resulted in a total computational cost of approximately 352,000 CPU hours. We make available the complete suite of data, encompassing both the raw CFD outputs and the derived post-processed datasets.
Our dataset serves as a benchmark for evaluating deep learning models, designed to facilitate effective model testing. To manage the large volumes of data from CFD simulations, we employ a data reduction strategy that focuses on key areas of the flow field. This involves retaining data from regions immediately in front of and behind the car, defined within a specific bounding box, which helps to reduce the overall data size significantly. Additionally, we provide a script that converts CFD simulation data into a format suitable for training deep learning models. Given the widespread use of data visualization tools such as ParaView and VisIt, which rely on the Visualization Toolkit (vtk), our dataset is made available in vtk format. This ensures that the data is easily accessible and usable within these common visualization environments, supporting a broad range of research and application needs.
The DrivAerNet dataset offers a comprehensive suite of aerodynamic data pertinent to car geometries, including key metrics such as the total moment coefficient , the total drag coefficient , total lift coefficient , front lift coefficient , and rear lift coefficient . Included within the dataset are crucial parameters like wall-shear stress, and the metric, integral for mesh quality evaluations. Further, the dataset provides insights into flow trajectories and detailed cross-sectional analyses of pressure and velocity fields along the and -axes, enriching the understanding of aerodynamic interactions.
The dataset includes:
Comprehensive CFD simulation data 16TB
A curated version of CFD simulations 1TB
3D meshes of 4000 car designs and the corresponding aerodynamic performance coefficients (, , , , and ) 84GB
2D slices include the car’s wake in the -direction and the symmetry plane in the -direction 12GB.
@@
Figure 7 displays three scatter plots that map out the relationships between drag coefficient () and various lift coefficients (, , and ) in the DrivAerNet dataset. The data has been split into training, validation, and test sets, with a division of 70 for training and 15 each for validation and testing. Such a division is critical for the integrity of the model training process and subsequent performance evaluation.
The Kernel Density Estimation (KDE) plot, illustrated in Figure 8, compares the distribution of drag coefficients across two aerodynamic datasets. Here, we compare the dataset from [36], which spans a broad range of drag values, reflecting a wide variety of car designs in ShapeNet. In contrast, our DrivAerNet dataset targets conventional car designs, accounting for more detailed geometric modifications. This focus is particularly relevant in the context of the engineering design process, where an initial car design is typically refined through incremental changes to optimize aerodynamic performance. The DrivAerNet dataset, therefore, provides a more specific examination of subtle design adjustments and their impact on aerodynamic performance.
In Figure 9, we present the aerodynamic performance across various designs. The top left illustrates the design with lowest drag coefficient . In contrast, the top right shows the design with the highest , identifying opportunities for aerodynamic refinement. The bottom left reveals the design with the lowest lift coefficient (indicating the largest downforce), which is advantageous for stability at high speeds, while the bottom right exposes design the highest , potentially complicating aerodynamic stability.

### v2

@@ 4 DrivAerNet Dataset
In this section, the workflow for generating the DrivAerNet dataset is presented in detail, as illustrated in Figure 1. The process begins with the selection of a baseline car model, chosen because it has comprehensive CFD and experimental data available. This baseline is critical for benchmarking and validating the accuracy of the simulations that follow. At this stage, a careful balance is struck between computational efficiency and simulation accuracy. Following this, a detailed parametric model is developed, enabling the generation of thousands of unique car designs. Automated meshing is then applied to these designs, preparing them for high-fidelity CFD simulations using OpenFOAM [20]. The resulting simulation data is systematically stored for subsequent analysis. Throughout the entire workflow, multiple quality control measures are integrated to ensure that the simulations not only converge but also maintain high accuracy.
@@
The goal is to create a large-scale dataset of 3D car shapes with high-quality mesh resolutions, watertight geometries, and diverse configurations that resemble real-world car designs. The ShapeNet [12] dataset offers a wide range of car shapes; however, the mesh resolution and non-watertightness make it unsuitable for large-scale engineering designs or CFD datasets. Other generic car models, like the Ahmed body [2] or the Windsor body [33], are very simple and do not represent actual cars with realistic shapes.
DrivAer model [23, 11] is a well-established conventional car reference model developed by researchers at the Technical University of Munich (TUM). It is a combination of the BMW series 3 and Audi A4 car designs in order to be representative of most conventional cars. The DrivAer model was developed to bridge the gap between open-source oversimplified models like the Ahmed [2] and SAE [13] bodies and the complex designs of the manufacturing companies, which are not publicly available. To accurately assess real-world aerodynamic designs, we selected the fastback configuration with a detailed underbody, wheels, and mirrors (FDwWwM) as our baseline model, as shown in Figure 2. This choice of the FDwWwM model was driven by the substantial impact of wheels, mirrors, and underbody geometry on aerodynamic drag, a conclusion supported by the findings in [22]. Specifically, the detailed underbody geometry adds 32-34 counts 111In aerodynamic evaluations, “drag counts” are used as a unit of measurement to denote small changes in the drag coefficient (). One drag count is defined as a increment in . to the drag, the inclusion of mirrors introduces an additional 14-16 counts, and the presence of wheels elevates the total drag coefficient by 102 counts.
@@
In order to create a comprehensive dataset for training deep learning models for surrogate modeling and design optimization, we first created a parametric model of the DrivAer model. This approach was necessitated by the limitations of the original model, which was provided as a single, non-parametric STL mesh. To adequately capture the geometric variations and design modifications relevant to real-world automotive challenges, we developed a version of the DrivAer model that is defined by 50 geometric parameters and includes 32 morphable entities (see Figure 2) using the commercial software ANSA®. This parametric model allows for a more detailed exploration of the design space, facilitating the generation of 4,000 unique design variants through the application of the Optimal Latin Hypercube sampling method, specifically employing the Enhanced Stochastic Evolutionary Algorithm (ESE) as outlined by [14].
@@
The parametric model, along with the constraints and the bounds applied during the Design of Experiments (DoE), significantly enriches the dataset, making it a robust foundation for the development and training of advanced deep learning models aimed at surrogate modeling and design optimization tasks. Diverging from the approach taken in [24], we implemented a broader range of morphing techniques, enabling us to explore a more diverse array of car designs. This approach aims to enhance the adaptability of deep learning models, allowing them to generalize across various car designs instead of being limited to minor geometric modifications within a single design. Figure 3 depicts the variation in mesh quality, ranging from coarse to high-resolution across various datasets. Our original mesh with 540k mesh faces, provides a denser and more detailed representation compared to the mesh resolutions in the studies by [28] and [44], thereby revealing more detailed geometric and design features.
@@ 4.1.1 Domain and Boundary Conditions
The DrivAer fastback model scaled 1:1 was selected for conducting the CFD simulations. The simulations were carried out using the open-source software OpenFOAM®, a comprehensive collection of C++ modules for tailoring solvers and utilities in CFD studies. In this study, the coupling between pressure and velocity was achieved through the SIMPLE algorithm (Semi-Implicit Method for Pressure Linked Equations), as implemented in the simpleFoam solver, which is designed for steady-state, turbulent, and incompressible flow simulations. The --SST model, based on Menter’s formulation [30], was chosen for the Reynolds-Averaged Navier-Stokes (RANS) simulations due to its ability to overcome the limitations of the standard - model, particularly its dependency on the freestream values of and , and its effectiveness in predicting flow separation.
The simulations were performed at a flow velocity () of 30 m/s, which corresponds to a Reynolds number of roughly , using the car length as the characteristic length scale. The computational mesh was constructed using the SnappyHexMesh (SHM) tool, featuring four distinct refinement zones. Additional layers have been added around the car body to precisely represent wake dynamics and boundary layer evolution (see Figure 5). Boundary conditions were defined with a uniform velocity at the inlet and pressure-based conditions at the outlet. To avoid backflow into the simulation domain, the velocity boundary condition at the outlet was configured as an outflow condition, with a specified inflow condition applied in cases of return flow. The car surface was assigned a no-slip boundary condition, while the wheels were modeled with a rotating wall boundary condition. Slip boundary conditions were applied to the lateral and top boundaries of the domain.
@@ 4.1.2 Validation of the Numerical Results
RANS fine | 205 | 0.81 | 1.88 |
The selection of the DrivAer fastback model is justified by the availability of both computational and experimental references, enabling us to benchmark our results against established data [22, 51]. Before commencing the simulations, we conducted a preliminary assessment of how mesh refinement influences the results. This involved comparing the drag coefficients obtained from two different mesh resolutions with experimental values and reference simulation, as detailed in Table 2. The objective was to identify an optimal balance between simulation accuracy and computational efficiency. This balance is crucial because our aim is to generate a large-scale dataset for training deep learning models, necessitating high fidelity in the simulation results, while also ensuring manageable disk storage and simulation times to accommodate the extensive computational requirements. The drag coefficient, , is determined by the equation:
@@
Our analysis reveals a consistent correlation between our simulations and both the benchmark experimental data and reference simulations. In computational fluid dynamics, especially in the preliminary stages of design, an error margin of up to 5 is generally considered acceptable for engineering purposes. Therefore, considering the balance between accuracy and computational resource allocation, we decided to conduct our simulations using the 8 million and 16 million cell meshes. These configurations provide a compromise between computational efficiency and the level of detail necessary for accurate aerodynamic analysis.
Rationale for Using RANS Simulations in Large-Scale Dataset Generation:
While the DrivAerNet dataset provides high fidelity in comparison to existing datasets, it remains limited in scope compared to the industrial standard, where simulations often involve transient hybrid RANS-LES or Wall-Modeled LES models and significantly larger computational grids (e.g., 100-200M cells). It is computationally prohibitive to run 4,000 simulations at this high level of fidelity. Therefore, in practical industrial applications, RANS simulations can be effectively employed to explore the design space during the initial or conceptual design stages. These simulations offer a balance between computational cost and accuracy, skimming the design space efficiently. In later stages of design, higher fidelity simulations, such as hybrid RANS-LES are typically used either for detailed CFD validation or for further training of machine learning models on more complex cases. As shown by [18], leveraging multi-fidelity CFD simulations is a robust strategy for accurate 3D flow field estimation. This approach combines the more easily obtainable RANS data, which captures general flow behavior, with high-fidelity Direct Numerical Simulation (DNS) data that provides detailed flow information despite being computationally expensive. Training a deep learning model with such a diverse dataset allows the model to generalize effectively to real-world scenarios, as confirmed by wind tunnel tests. This two-phase training strategy begins by using medium-fidelity RANS data to capture general flow patterns and subsequently fine-tuning the model with DNS data to enhance both accuracy and real-world applicability.
Similar results leveraging multi-fidelity datasets for surrogate model training have been demonstrated by [41, 43], underscoring the efficacy of this approach in aerodynamic analysis. The DrivAerNet dataset can also be utilized in this manner, allowing integration with datasets of either lower or higher fidelity to augment model training and refine predictive capabilities.
@@ 4.1.3 CFD Simulation Results
In contrast to the approach by [44] where all car models are standardized to a uniform length of 3.5 meters to fit a predefined computational domain, our dataset allows diversity in car dimensions, adjusting the mesh, bounding boxes, and additional layers accordingly for each design. This flexibility is crucial for capturing the intricate flow dynamics around cars, including phenomena like flow separation, reattachment, and recirculation, and ensuring precise aerodynamic coefficient estimation. This approach addresses limitations observed in some studies that prioritize dataset size over simulation fidelity, often overlooking the importance of convergence, accurate modeling, and appropriate boundary conditions for complex 3D models.
@@ 4.2 Geometric Feasibility
In our approach to generating large diverse automotive designs through automated morphing with ANSA®, ensuring the geometric quality and feasibility of each variant is crucial. To address potential issues such as non-watertight geometries, surface intersections, or internal holes resulting from the morphing operations, we employ an automated mesh quality assessment and repair process. This procedure not only identifies but also rectifies common geometric anomalies, ensuring that only simulation-ready models are included in our dataset. Geometries failing to meet these criteria are systematically excluded from simulations. The DrivAerNet dataset employs a diverse array of parameters (total of 50 parameters) to morph car geometries, encompassing modifications to the side mirror placement, muffler position, windscreen, rear window length/inclination, the size of the engine undercover, offsets for the doors and fenders, hood positioning, and the scale of headlights as well as alterations to the overall car length and width. Additionally, adjustments are made to the car’s upper and underbody scaling, as well as pivotal angles such as the ramp, diffusor, and trunk lid angles, all crucial for exploring the effects of different design modifications on car aerodynamics. For a detailed account of the morphing parameters, including their lower and upper bounds, please refer to our GitHub repository.
As we morph the entire car’s geometry, wheel positioning is adjusted in the , , and axes during the morphing process. For all simulations, we use front and rear wheels of the same shape. To accurately simulate the wheel rotation, we export them as separate STL files post-morphing, which allows us to apply a rotating wall boundary condition. Moreover, morphing affects the car’s vertical positioning, necessitating a calculation of the -axis displacement to ensure the car body and wheels are properly aligned with the ground plane. For simulation purposes, we supply three distinct STLs to model their interactions accurately: one for the car body, one for the front wheels, and one for the rear wheels.
@@ 4.3 DrivAerNet Dataset Characteristics
For our simulations, we employed OpenFOAM® version 11 [20], executing the computational tasks across 128 CPU cores and 4 Nvidia A100 80GB GPUs. This resulted in a total computational cost of approximately 352,000 CPU hours. We make available the complete suite of data, encompassing both the raw CFD outputs and the derived post-processed datasets.
Our dataset serves as a benchmark for evaluating deep learning models, designed to facilitate effective model testing. To manage the large volumes of data from CFD simulations, we addressed dataset size and accessibility by dividing DrivAerNet into smaller subsets tailored to specific tasks, reducing storage needs. For example, the 4000 car geometries in STL format occupy 84 GB, while aerodynamic performance coefficients, including the total moment coefficient , total drag coefficient , total lift coefficient , front lift coefficient , and rear lift coefficient , are stored in CSV format. The full CFD volumetric field, which contains velocity, pressure, and wall-shear stress data, occupies 5.6 TB in VTK format. Surface pressure field data in VTK format totals 102 GB, and wall shear stress (WSS) data sums up to 200 GB, also in VTK format. We have also partnered with Harvard Dataverse to provide easy access to these subsets, allowing researchers to download only the data they need, thereby minimizing storage requirements. A GitHub repository will be updated with detailed subset information and download guidelines. Given the widespread use of data visualization tools such as ParaView and VisIt, which rely on the Visualization Toolkit (VTK), our dataset is made available in VTK format. This ensures that the data is easily accessible and usable within these common visualization environments, supporting a broad range of research and application needs.
The DrivAerNet dataset features full 3D flow field information, as shown in Figure 6(a) with the velocity data, and in addition, it provides the pressure distribution on the car’s surface. The pressure coefficient, denoted as , is calculated as the ratio of the pressure differential to the dynamic pressure, , and is given by:
The distribution of on the car surface is depicted in Figure 6(b).
The DrivAerNet dataset offers a comprehensive suite of aerodynamic data pertinent to car geometries, including key metrics such as the total moment coefficient , the total drag coefficient , total lift coefficient , front lift coefficient , and rear lift coefficient . Included within the dataset are crucial parameters like wall-shear stress and the metric, integral for mesh quality evaluations. Further, the dataset provides insights into flow trajectories and detailed cross-sectional analyses of pressure and velocity fields along the and -axes, enriching the understanding of aerodynamic interactions.
An overview of the dataset components can be found in Table 3, which details the types, formats, and information provided for each dataset element, including the 3D car geometries, aerodynamic performance coefficients, full CFD volumetric fields, and surface fields like pressure and wall-shear stresses.
Dataset Component | Format/Type | Information Provided |
3D Car Geometries | STL | Surface representation of the car geometries |
Aerodynamic Performance Coefficients | CSV | Total moment coefficient , total drag coefficient , total lift coefficient , front lift coefficient , and rear lift coefficient |
Full CFD Volumetric Field | VTK | Velocity field, pressure field, wall-shear stresses |
Surface Pressure Field | VTK | Pressure distribution on the car surface |
Wall Shear Stresses | VTK | Wall-shear stresses on the car surface |
@@
In Figure 7, we present the aerodynamic performance across various designs. The top left illustrates the design with the lowest drag coefficient . In contrast, the top right shows the design with the highest , identifying opportunities for aerodynamic refinement. The bottom left reveals the design with the lowest lift coefficient (indicating the largest downforce), which is advantageous for stability at high speeds, while the bottom right exposes the design with the highest , potentially complicating aerodynamic stability.
Figure 8 illustrates scatter and kernel density estimation plots showing the relationships between the drag coefficient () and lift coefficients () for the DrivAerNet dataset. The data points represent unique design variants generated via Optimal Latin Hypercube sampling and the Enhanced Stochastic Evolutionary Algorithm (ESE). The dataset is categorized into training, validation, and test sets, with a division of 70 for training and 15 each for validation and testing. This division is critical for maintaining the integrity of the model training process and ensuring robust performance evaluation.
Analysis of Datasets Diversity Using Chamfer Distance
The Chamfer distance is a metric used to quantify the geometric diversity between two point sets, making it highly applicable for analyzing the diversity in datasets of 3D models [19]. It is defined as follows:
Where:
represents the squared Euclidean distance between point from set and point from set .
is the minimum squared distance from any point in to points in .
is the minimum squared distance from any point in to points in .
This measure allows for the calculation of a diversity score by averaging pairwise Chamfer distances across all model pairs within a dataset, offering a quantitative measure of model dissimilarity. We computed the diversity scores for both the Song et al. [44] dataset (based on ShapeNet) and DrivAerNet datasets. ShapeNet achieved a diversity score () of 0.433, demonstrating a wide range of car shapes including trucks, sports cars, and various utility cars, which introduces substantial shape variation. Conversely, DrivAerNet recorded a diversity score of 0.171, concentrating on more precise geometries tailored to conventional car designs. This targeted focus aids in enhancing model performance by enabling detailed learning from subtle geometric variations within a more restricted range of automotive shapes.
This concept of diversity is crucial when examining the practical impacts of these variations on aerodynamic efficiency, as depicted in the Kernel Density Estimation (KDE) plot shown in Figure 9. The KDE plot compares the distribution of drag coefficients between the broad-ranging ShapeNet dataset and our more specialized DrivAerNet dataset. While ShapeNet covers a broad spectrum of drag values reflecting its diverse car designs, DrivAerNet focuses on conventional car designs with detailed geometric modifications, pertinent to the engineering design process. In this process, an initial car design is incrementally refined to optimize aerodynamic performance. Therefore, the DrivAerNet dataset provides an in-depth look at how subtle design changes affect aerodynamic performance, underscoring the practical application of diversity scoring in real-world engineering scenarios. While ShapeNet offers design diversity, its poor mesh quality and geometry issues make it unsuitable for high-fidelity CFD simulations. In contrast, DrivAerNet’s parametric design ensures high-quality, automated meshing, leading to more accurate and reliable CFD results, even with a more focused range of designs.